Observing Run 2008-I may 17th to june 5th
June 2nd: Observations (Karine,
- Relatively bad night, r0 lower than 5 cm, wind
- * Validation of secondary fringe tracker
- * Series of polarized and natural data for
calibration of the SPIN mode on two stars but low quality data
- * Observations of gLyr/bLyr/gLyr in medium
resolution near Halpha but very low SNR due to seeing
- * Attempt in the low resolution mode on
HD181655 (mV = 6.3) with its calibrator.
June 1st: Observations
(Karine, Denis) (logbook)
- Relatively good night, r0 around 7cm (5-12 cm)
- Validation of baseline switching. 2 simultaneous baselines
used and we switch from one to the other just by using the shutters. It
works pretty well in less than 5 mn. Observations on delta and gamma
Uma for calibration purposes on S1S2 and W1W2. The only remaining
problem at this stage is that we did not succeed to speak to the second
OPLE server, excepted manually on ctrscrut.
- Program Be: Medium Resolution, delta Sco + 1 calibrator.
- Program rotator: High resolution, alpha Cep (Halpha) at
-2h05, -1h50, -1h00, -0h25 and (Fe528nm) at HA=-1h30 and -1h10
- Program: Medium resolution P Cygni + ups Cygni and 2
S1S2, W1W2. On Cal2 (mV=4.9, DeltaLambda=50nm), fringe tracking Ok,
fringes very stable and SNR about 20 in few tens of seconds.
- Just a small note: on W1W2, we noticed large errors (0.5 to
1.2 µm) on
W2-OPLE position (W2 was the reference cart) and clearly the fringes
were blurred on our detector...strange...
- Otherwise, every thing was really great!!!!!
- Cerise sur le gateau...(in french in the text), the last
observing sequence on alpha Cep was partially controlled by our friends
in Grasse. Every things appear to work properly excepted the dialog
with the OPLE. Next try in the coming days.
May 31th: Observations
(Karine, Denis) (logbook)
- Good night, seeing between 8 and 12cm, very stable. Small
- S1 on beam 5 and S2 on beam 6 for finding the fringes with
CLASSIC following our last night failure. Fringes were expected to be
around -3mm but were found at +27.4mm!
- S1 on beam 3, S2 on beam 2 for VEGA. Fringes found at
Nice recording of very high SNR fringes.
- on delta UMa, we test the changes of beams, S1 on beam 2
on beam 1. It tooks about 10mn for changing the configuration and the
fringes are found at 27.9mm. New recording with very good data.
- After UT=6h00 science programs on S1 on beam 3 and S2 on
- Rotator: Alpha Cep (high spectral resolution). First around
Halpha and then Hbeta. On both configurations we succeeded to close the
loop for the fringe tracking. Nice data recorded from HA=-5h50 to
- Be: delta Sco + 2 calibrators. Medium spectral resolution,
- Rotator: delta Aql, Medium resolution: 2 observations +
calibrators (HA=-1h, HA=+1h) + one point with the high resolution
(mV=3.4, R=30000! and fringe tracking).
- Every thing works perfectly tonight. It was really a great
(after so many years...) to see the very nice stability of stars and
fringes, the fringe tracking working well and all these data!
May 30th: Observations (Karine,
- Low RH but seeing variable this night: from 5 to 10cm in
the first half
then around 12cm but variable.
- alfCVn + del UMa on W1W2 rapidly for qualification of
beams3&4. Polarization impossible on beam 4 (vignetting on the
- Look for fringes on S1S2 but not found from -12.5 to
Seeing not very good but we should have seen them. Strange...Metrology
- Again on W1W2: long serie on zet Oph (-1h15->+2h), ups
(-2h20->0h0) with calibrators: HD170296, HD193432, HD198001 but
fringes very low on the first two calibrators and two sequences on the
last one. For ups Sgr, we were surprised to see the quite high
visibility wrt what we expected.
- Good coherence tracking on mV=4.6, V² around 0.35 and
- Too short at the end to have a calibrated point on bet Lyr.
May 29th: Observations (Karine, Denis) (logbook)
May 28th: Observations (Karine,
- Very good night but early closing due to RH. Seeing around
12 to 15cm.
- ~UT4:00: Final alignment of CHARA beams on the VEGA
positions by moving the TT mirrors for the images and the M10 (or POP
mirror) for the pupils. It works very well and images and pupils could
be correctly centered. Our first target was difficult in fact and
fringes were difficult to find. Bad choice...
- ~UT5:30: Validation of the group delay tracking with VEGA
time processing. Corrections of typically a few µm are sent
automatically to the movable cart to block the OPD at a fixed position
within the envelope of coherence. Long integrations (of spectral
densities) are thus possible.
- ~UT6:00: Data recorded on delta UMa in natural light and
polarized light for qualification purposes.
- UT6:45->UT10:40: Long sequence with calibrators around
lyrae: eps Her, del Her, eps Her, bet Lyr, gam Lyr, bet Lyr, gam Lyr,
bet Lyr, del Her (bad), eps Her (bad).
- Standby due to humidity no-reopening possible! Too bad...
- The control system of VEGA is now well stabilized and
is very easy!
- Late opening due to high RH. 4 hours of observations.
Fringes ok on W1W2 but important pbs of metrology.
- Validation of the Vega Group Delay Tracking
May 27th: Changing the
crew..+ first observation (..) (Karine, Denis)
- Jean-Michel and Aurélie are now back to France and
Karine is arrived.
- First clear night but humidity drops down only around
midnight. Then a few problems with CHARA due to the power outage of
this week-end. No fringes on VEGA...no fringes on CLASSIC: POP
configurations was badly read by GrandWazoo! After many attemps we
succeeded to find the fringes on CLASSIC but it was about half an hour
before sunrise. No fringes found on VEGA. Hopefully next night!
May 26th: Bad conditions.
Software (Aurélie, Jean-Michel, Denis)
- The bad conditions are always here. Fog and humidty!
- Installation of the new control computer (difficult...) and
various upgrades at different levels
- No observations.
May 24th, May 25th:
Software, Calibrations. (Aurélie, Jean-Michel, Denis)
May 23th: Snow and
- The bad conditions are continuing. Fog, humidity, cold
temperatures. These two days have been used to work on the control
software, the data archiving facility, the detector software and the
group delay tracking software. We have also redo the determination of
the spectral laws and recorded new spectra with the new detector's
software. We have also written a certain number of documentations for
- Every thing seems ready for observations.
- Snowy wake-up this morning! The outside temperature started
at 32°C 4 days ago and goes down to 2°C today!!!
- Long, long, long work on the detectors but great successes.
They are now under control. For more details see report (here,
sorry in french). The main problem
was that structures on the background of the detectors were introducing
great artefacts in the data, in the spectrum and in the spectral
densities. These features are now completely removed. The problem
originates in the readout mode of the CMOS detector where the parameter
(frame rate and exposure time) were not correctly set. Now we are
working with a frame rate of 40 fps and an exposure time of 15ms.
- Finally we obtained a correct calibration of the
spectrograph in medium resolution (see doc)
with lambda=650nm in the red camera, giving lambda=500nm in the blue
one. This will allow simultaneous recording of Halpha and Hbeta.
- Any possible observations yesterday and tonight...Too bad!
May 22th: Wind, hail,
hamburger, outlet...and detectors (Aurélie, Jean-Michel, Denis)
May 21th: Detector,
(Aurélie, Jean-Michel, Denis)
- No observations yesterday due to high winds and lot of
- Today, really bad weather with hail in the afternoon and
high humidity preventing for observations
- The software configuration of the DALSA cameras seems to
contain some problems perturbating the signal amplification of the
signal in the pixels. After a lot of tests on the configuration of the
various parameters of the configuration we identified some troubles
coming from potential conflicts between the applications. We finally
to obtain a situation where the images are really very good without
perturbations column by column but the situation is not completely
stable at that time. More tests are necessary.
- Today, early start due to the huge wind near the Monastery!
- Almost all the day has been spent to work on the detector
software. We have tested the different versions of the program
containing or not the real time correction of the flat field of the
detector before photon detection. The main residual problem is the
fixed pattern on the integration of the images, leading to a very
important noise on the spectrum. Hopefully due to the important motion
of fringes and speckles on the detector, we do not expect an important
effect on the visibilities. This pattern did not disappear even with a
flat field correction before the photon detection. The reason is that
we use a high gain and skimmer gain on these DALSA cameras allowing a
high sensitivity but a large noise on the dark! Thus the flat field
correction is not stable and the pattern on the integration of the
images did not diminish with the integration time. The conclusion is
that we will continue to use the current version without the flat field
correction with a more strict filtering of the very low level photons.
- In parallel we have been able to fix the correct dispersion
law in the medium spectral resolution. On the red camera around 656nm,
the spectral band is 49nm, whereas it is 35nm in the blue camera around
500nm. The spectral distance between the two cameras is 153nm. Thus the
simultaneous observation in Halpha and Hbeta is possible by setting the
red camera around 645nm.
- Great progresses also on the VEGA documentation (alignment,
- Tonight the plans are to test the detector software on the
sky, the fringe tracking operation and the observing scheduler. However
high winds and lot of dusts prevent us to open at that time...Forecast
for the coming days are not very good with clouds and thunderstorms.
May 20th: Cophasing and
(Aurélie, Jean-Michel, Denis)
- A translation of x mm of the BCP introduces a 4x mm OPD for
the retro beams and not a 2xmm OPD! So the BCP have to be shifted by
half the measurements. We finally succeed to fully cophase (within
50µm) the four VEGA beams with respect to the CHARA beams. BCP#4
has been translated by 0.7mm, BCP#2 by 0.85mm and BCP#1 by 3.35mm. The
alignment has been corrected a little bit to compensate the angular
errors introduced during the transtalion. Things are now stabilized and
it looks pretty nice!
- Laszlo has fixed the periscope issue.
- A lot of time has been spent on installing the new central
control computer. Not yet finished...linux...
- We started intensive tests of the new version of the
control software. Bugs are corrected and things are progressing.
- Yesterday night we spend a few hours on the night on the
visible beams of the FLUOR programs. However problems with CHARA
software prevent us to get fringes during the first half of the night.
May 19th: Cophasing and
(Aurélie, Jean-Michel, Denis)
- Thanks to Judit, the CHARA internal fringes were acquired
on beams 3&2 on the CHARA spectrograph and then the metrological
corner cubes have been correcly installed. The VEGA white light source
is used back to CHARA and then back to VEGA. Thanks to small astuces to
correctly separate the direct beams to the spectrograph from the beams
coming to the spectrograph after going to the CHARA corner cubes we
succeeded to detect the retro beams from our white light source in the
high spectral resolution mode of the spectrograph. Fringes on beams
2&3 were found with an offset of 1.7mm and we deduce that BCP#2 has
to be shifted forward by 1.7mm.
- Second step was to go on beams 3&4 but the beam coming
back were not correctly superimposed. The only solution we have found
is to replace the cornercubes by tiltable flat mirrors. Then the beams
were correctly superimposed on the spectrograph slit and the fringes
were found on beams 3&4 with an offset of 1.4mm.
- Due to the preparation of the night, we just have had time
to correctly translate BCP#4 by 1.4mm and to control that the
alignement has been kept correct. Next step is to verify that the
fringes are centered and then that VEGA beams 3&4 are cophased wrt
CHARA. Then translation of BCP#2, verification, measurements of fringes
on beams 1&2 and translation of BCP#1 and verification.
- Various software update (windows, compiler, automatic ntp
- Setting of the data archiving facility: installation of
removable hard drives on the data reduction
computer for the storage of the raw data. All the data are transferred
from the various computers through network NFS mounting. Every thing is
execute an archive script at the end of each observing night. At the
end of each run, a pair of hard drives is returning to France for the
scientific data analysis.
May 18th: Start of the
technical part of the run (Aurélie, Jean-Michel, Denis)
- We found the instrument perfectly aligned since last
december. Due to recent actions on the CHARA periscope mirrors, we have
control and finely tune the VEGA-CHARA Alignment. We are now ready for
the cophasing operation. The trim angle of the grating turret has been
corrected, as identified on the last data. The detectors have been
cleaned and their orientation corrected. The new filter wheel in front
of the Alignment Unit is in operation. We have also recorded important
data files for the data pipeline such as the images of the internal
pupils and images as well as raw data files of optically simulated
- The hardware parts of the new control system have been
installed and tested. This concerns mainly the internet and electrical
connections of the different subsystems for a full remote control. The
new versions of the control software are progressively installed and